130-J Remington Ave., Syracuse, N.Y. 13210

November 8, 1999

Easter Service Center, Immigration and Naturalization Service, 74 Lower Weldon St. St. Albans, Vt 05479-9765

Dear Sir:

This letter is written to request replacement of my lost employment authorization document as a J-2 Dependent. My husband is a duly enrolled full-time student at Syracuse University. He is supported by a scholarship from the Northeast Parallel Architecture Center(NPAC) at Syracuse University. He receives $1000 per month. These funds plus savings are sufficient to support him and his dependent.

My application for permission to work is based on my desire to remain active, gain experience I can use when I return to my home country, and to provide funds cultural, recreational, and travel expenses associated with experiencing American culture and customs. I understand the limitations of J-2 employement. My income will be used only to support the activities listed above, and will not be used to support my husband.

Appropriate documents for this request are enclosed. Thank you for your assistance in granting this employment authorization.

Sincerely,

Zeynep (Odcikin) OZDEMIR